How to do this sum: 2 over x-y plus 1 over x( to the power of two) - y( to the power of two)
1
Expert's answer
2013-02-15T04:14:46-0500
We have the following statement:

A = 2/(x-y) + 1/(x²-y²).

First, let's reduce this statement to the common denominator. The common denominator is

(x²-y²),

as

(x²-y²) = (x+y)(x-y).

Therefore

A = 2/(x-y) + 1/(x²-y²) =

= 2(x+y)/((x-y)(x+y)) + 1/(x²-y²) =

= 2(x+y)/(x²-y²) + 1/(x²-y²) =

= (2(x+y) + 1)/(x²-y²) =

= (2x+2y+1)/(x²-y²).

So,

2/(x-y) + 1/(x²-y²) = (2x+2y+1)/(x²-y²).